Suppose that a lender's desired real rate of interest is 2%, the expected rate of inflation is 2% and the actual rate of inflation is 4%: a) What's the nominal rate of interest?

b. What's the actual rate real rate of interest? Explain who benefits from expectations error--the lender or the borrower.
Answer #1

a)

By approximation formula

Nominal interest rate=desired real rate of interest+expected inflation rate=2%+2%=4%

By exact formula

Nominal interest rate=(1+desired real rate of interest)*(1+expected inflation rate)-1

=(1+2%)*(1+2%)-1=4.04%

b)

By approximation formula

Actual real rate of interest=Nominal rate of interest-Actual inflation rate=4%-4%=0%

By exact formula

Actual real rate of interest=(1+Nominal rate of interest)/(1+Actual inflation rate)

=(1+4%)/(1+4%)-1=0%

Clearly borrower is paying less real rate of interest because of expectation error in this case.So, borrower will be benefited.